Melissa's brother is 6 years less than twice her age. The sum of her age and her brother's age is 27. Which equation best shows this information?

Found 2 solutions by ankor@dixie-net.com, Quadratic1600:
Answer by ankor@dixie-net.com(22740)

Melissa's brother is 6 years less than twice her age.
b = 2m - 6
The sum of her age and her brother's age is 27.
b + m = 27
Which equation best shows this information?

Answer by Quadratic1600(28)

Hi
The best way and possibly the only way to work out this question is to create 2 simultaneous equations based on the given information.
So lets have a look
Lets say Melissa is M and her brother is B
So
B = 2M - 6 (A)
M + B = 27 (B)
Ok so lets use the substitution method and put (A) in to (B)
Since B = 2m - 6 then we can say
M + 2M - 6 = 27
3M - 6 = 27
3M = 33
M = 11
Now M + B = 27 so 11 + B = 27
B = 27 - 11
B = 16
Lets check that we are right
Melissa's brother is twice her age which would be 22 and then we minus the 6 which would be 16 so we are correct.
I hope you understand the main principle here but if not by all means send a message and I will go through it again.